# Anisotropic resistivity of a $p$-wave magnet candidate CeNiAsO

### v1

@@ Abstract
S. L. acknowledges helpful discussions with Prof. Quansheng Wu and Prof. Guoqiang Yu. This work is supported by the National Key Research and Development Program of China (Grants No. 2022YFA1403400, No. 2021YFA1400400, No. 2022YFA1403800), the National Natural Science Foundation of China (Grants Nos. 12274411, 12241405, and 52250418), the CAS Superconducting Research Project (Grant No. SCZX-0101), the Basic Research Program of the Chinese Academy of Sciences Based on Major Scientific Infrastructures (Grant No. JZHKYPT-2021-08), and the CAS Project for Young Scientists in Basic Research (Grant No. YSBR-084). This study was also partially supported by Synergetic Extreme Condition User Facility (SECUF).

### v2

@@ Abstract
S. L. acknowledges helpful discussions with Prof. Quansheng Wu and Prof. Guoqiang Yu. This work is supported by the National Key Research and Development Program of China (Grants No. 2022YFA1403400, No. 2021YFA1400400, No. 2022YFA1403800), the National Natural Science Foundation of China (Grants Nos. 12274411, 12241405, 52250418, and 12374143), the Basic Research Program of the Chinese Academy of Sciences Based on Major Scientific Infrastructures (Grant No. JZHKYPT-2021-08), and the CAS Project for Young Scientists in Basic Research (Grant No. YSBR-084). This study was also partially supported by Synergetic Extreme Condition User Facility (SECUF).